# Speech Robust Bench: A Robustness Benchmark For Speech Recognition

### v2

@@ Abstract
As Automatic Speech Recognition (ASR) models become ever more pervasive, it is important to ensure that they make reliable predictions under corruptions present in the physical and digital world. We propose Speech Robust Bench (SRB), a comprehensive benchmark for evaluating the robustness of ASR models to diverse corruptions. SRB is composed of 114 input perturbations which simulate an heterogeneous range of corruptions that ASR models may encounter when deployed in the wild. We use SRB to evaluate the robustness of several state-of-the-art ASR models and observe that model size and certain modeling choices such as the use of discrete representations, or self-training appear to be conducive to robustness. We extend this analysis to measure the robustness of ASR models on data from various demographic subgroups, namely English and Spanish speakers, and males and females. Our results revealed noticeable disparities in the model’s robustness across subgroups. We believe that SRB will significantly facilitate future research towards robust ASR models, by making it easier to conduct comprehensive and comparable robustness evaluations.
@@ 1 Introduction
As novel ML models continue to be developed and deployed at an ever-increasing rate, it has become crucial to ensure that their robustness to corruptions arising from a myriad of sources, including the environment, sensing apparatus, and even malicious actors is well understood. To this end, prior works have developed comprehensive robustness benchmarks, particularly for vision [1, 2, 3, 4] and natural language processing models [5, 6], that evaluate several aspects of a model’s performance under a variety of input perturbations. These benchmarks have proven to be invaluable to the advancement of research into more robust models because (1) they unify robustness evaluations, thus enabling meaningful comparisons across models and allowing progress to be accurately tracked, and (2) they make it easier for researchers to comprehensively evaluate the robustness of their models by aggregating a diverse and representative set of perturbations in a single benchmark.
@@
While several robustness benchmark datasets exist for Automatic Speech Recognition (ASR) models [7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18], none of the currently existing ones is in any sense comprehensive: each benchmark measures the model’s robustness w.r.t. to one or a few specific types of perturbations, thereby putting the onus on model developers to find and collect all the relevant benchmarks to evaluate their model comprehensively. This has often resulted in model developers evaluating their models on disparate benchmarks  [19, 20, 21, 22], which makes it hard to reliably compare performance and robustness across models. Recently, Huggingface Open ASR Leaderboard [23] has sought to unify ASR model evaluations by developing a benchmark consisting of several real-world speech datasets. Although evaluating models on exclusively natural data may accurately reflect their real-world performance, it is generally not informative about the specific types of perturbations the models are weak or strong against, because the noise sources present in these datasets are not controlled or even fully known. For example, the crowdsourced recordings in Common Voice [24] contain a variety of distortions including sensor noise from low-quality equipment, background noise, and mispronunciation by non-native speakers. Additionally, crucial digital perturbations like special effects, computer-generated speech, and adversarial examples, to which humans are usually invariant, are largely overlooked by existing benchmarks.
@@
In this paper, we propose Speech Robust Bench (SRB), a benchmark for comprehensively evaluating the robustness of ASR models to input perturbations and corruptions. SRB is designed to address the aforementioned major shortcomings of existing ASR robustness benchmarks, i.e., that (1) they are often specialized and thus are not individually comprehensive, (2) even taken together, they overlook important perturbation types. SRB addresses these shortcomings by curating a large comprehensive bank of 114 perturbations and inherently noisy data sources. The perturbation bank includes distortions and variations caused by the environment, recording equipment, special effects, computer-generated speech, and significantly, adversarial attacks, which are often overlooked by existing benchmarks. When selecting inherently noisy data sources, we take special care to ensure that the sources of variation in each data source are limited and known, so that the models’ specific weaknesses can be identified. These sources contain speech with non-native English accents and inter-personal conversations, giving rise to articulatory and lexical variability in speech.
@@
To highlight the need for and the benefits of doing systematic and comprehensive robustness assessment, we evaluate the robustness of several popular ASR models (§4.1) using SRB. We observe that Whisper [19] is the most robust on average among the models we tested, even outperforming the more recent Canary [25]. We conduct further analyses to disentangle the effects of model and training data size, revealing that larger models tend to be more robust than smaller models, even if the latter are trained on significantly more data. We further extend our analysis by evaluating the models’ robustness for the various population sub-groups, namely, English and non-English (Spanish) speakers, and male and female speakers. We find that significant disparities exist across these sub-groups, thus identifying issues with fairness of the models and highlighting areas where future work could provide clear improvements. Besides pinpointing robustness issues, this demonstrates the utility of SRB for fairness evaluations that consider robustness disparities across models as well.
@@
  • To facilitate out-of-the-box robustness evaluations for the research community, we have publicly released a large dataset (https://huggingface.co/datasets/mshah1/speech_robust_bench) containing perturbed versions of LibriSpeech [26], TEDLIUM [27] and Spanish Multilingual LibriSpeech [28] test sets, and accented speech from common voice and segemented near- and far-field audios from CHiME-6 [10] and AMI[8].
  • We open source our code with clear documentation to allow for easy reproducibility and extensibility.(https://github.com/ahmedshah1494/speech_robust_bench)
@@ 2.1 Robust Automatic Speech Recognition
Several techniques have been proposed for making Automatic Speech Recognition (ASR) models robust to input perturbations, such as noise and other signal corruptions [29]. We can divide these techniques into two high-level categories: i) model-based and ii) feature-based. Model-based techniques modify the models to make them more robust. Examples of such approaches include adapting pre-trained models [30, 31], de-noising the audio before processing [32, 33], and training ASR models on noisy data [22]. Since model-based strategies generally require access to noisy data [29], they are most effective if the sources of noise, and/or the exact environment in which the ASR model will be deployed in are known, and one can gather data to represent them. Feature-based approaches, on the other hand, involve developing handcrafted features, invariant to noise and corruptions in the signal [29]. Several of these features are inspired by biological audition [34, 35, 36], while others use signal processing techniques [29]. Generally, these methods are designed to extract the components of the audio signal salient for speech production and perception, while discarding irrelevant components [37]. Consequently, they do not require precise knowledge of the environment and noise distributions. Recently, however, handcrafted features have fallen out of favor, and have been replaced by features learned via end-to-end training of deep learning models on large amounts of data [38, 39, 22, 19]. Proponents of these techniques posit that models trained on larger datasets become more robust. Our evaluations in § 4 reveal that there are several input perturbations against which smaller models trained on less data outperform larger models trained on more data.
@@ 2.2 Adversarial Robustness
Adversarial perturbations can change the response of a model when added to their inputs, but are either imperceptible to humans or perceptually and semantically irrelevant enough to be ignored by them [40, 41]. Adversarially perturbed inputs are known as adversarial attacks. They can be targeted (aiming to change a prediction to a specific incorrect class), or un-targeted (aiming to change a prediction to any incorrect class [42]). The design of adversarial attacks is determined by the level of knowledge the attacker is assumed to have about the target model. Attacks that assume full knowledge of the target model’s architecture and weights (white-box threat model) often use gradient-based optimization techniques [40, 41, 43, 44, 42]. Attackers who do not have any knowledge of the target model’s architecture and only have query access to it (black-box threat model) typically use gradient-free optimization methods [45, 46, 47, 48, 49, 50]. An intriguing property of adversarial perturbations is that they transfer between models [51], and inputs [42, 52], i.e. perturbations designed for one model/input may be effective against others as well. Our SRB includes two types of white box adversarial attacks; those that generate perturbations: 1) specific to each input [43], 2) that cause models to mis-transcribe multiple inputs [52].
@@ 2.3 Robustness Benchmarks for Speech
Several robustness benchmark datasets exist for Automatic Speech Recognition (ASR) models [7, 8, 9, 10, 11, 12, 9, 13, 14, 15, 16, 17, 18, 53, 54, 55]. Unfortunately, these benchmarks do not comprehensively evaluate model robustness. While past works such as [54] did indeed propose benchmarks containing diverse perturbations, many of the recently proposed benchmarks measure the model’s robustness to one or a few specific types of perturbations. For example, some datasets [15, 17, 18, 16] focus exclusively on reverberant speech, while others focus only on multi-speaker scenarios [8, 7], environmental noise [14, 10, 9, 53], or accented speech [56, 57]. In this situation, the onus is on model developers to collect all relevant benchmarks and comprehensively evaluate their model. Furthermore, certain types of digital perturbations, like special effects, computer-generated speech, and adversarial examples, which humans are usually invariant to (and thus ASR models are expected to be as well) are largely overlooked by existing benchmarks. To facilitate future research, SRB combines several benchmark datasets containing known noise sources to allow researchers to comprehensively evaluate the robustness of their models using a single benchmark.
@@
In the absence of comprehensive benchmarks, prior works have tried to evaluate ASR model robustness in various ways. For example, several works have used subsets of existing benchmarks to evaluate the model robustness [19, 20, 21], while others [22, 19, 58] have evaluated the robustness of models by computing their transcription accuracy on multiple natural speech datasets. A key shortcoming of these approaches is that the choice of benchmarks and datasets varies across studies, making it difficult to compare results. Recently, Huggingface Open ASR Leaderboard [23] has sought to mitigate this issue of inconsistent evaluation setups by developing a benchmark comprising several real-world speech datasets. Although evaluating models only on naturally noisy data may accurately reflect their real-world performance, it may not inform about the specific type of perturbations the models are weak against, because noise sources present in these datasets are not controlled or even fully known. For example, the crowdsourced recordings in Common Voice [24] contain a variety of distortions including sensor noise from low-quality equipment, background noise, and mispronunciation by non-native speakers. If a model struggles with Common Voice, identifying the specific types of noise it is sensitive to is not a straightforward task.
By evaluating models on a diverse set of known perturbations and noise sources, including adversarial attacks, SRB complements and extends benchmarks, like [23], that measure average case real-world performance and, thus, allows the robustness of ASR models to be comprehensively evaluated.
@@ 3 Speech Robust Bench
At a high level, Speech Robust Bench (SRB) consists of four components: (1) Clean datasets, and a curated selection of inherently noisy datasets containing limited and known sources of noises and variations that are difficult to simulate; (2) Large bank of perturbations applicable on audio data; (3) Transcription extraction from ASR models and analysis; (4) Set of metrics for quantifying ASR model robustness, and facilitating comparisons among models. The basic workflow of SRB is the following (see Fig. 1): first, recordings from the clean dataset(s) (§ 3.1) are perturbed by a large bank of perturbations (§ 3.2), resulting in several perturbed versions of each recording. Next, these perturbed recordings, the original clean recordings, and the recordings with inherent noise are transcribed using the target ASR model. The predicted and reference transcripts are compared, and the accuracy and robustness of each model under each type of noise are captured with various metrics (§ 3.3). To account for the relative differences in the level of distortion introduced by different perturbations, we estimate speech quality scores using appropriate models and use them to normalize the metrics.
3.1 Dataset
We use clean speech from Librispeech [26] and TEDLIUM [27]. Librispeech contains professional recordings of audio books, and thus, has properly enunciated and cleanly recorded speech. Meanwhile, TEDLIUM contains professional recordings of TED talks and provides a degree of lexical and phonetic diversity that may be lacking in Librispeech. We chose these datasets due to their popularity in the ASR literature, however, any dataset containing utterances and reference transcripts would also be compatible with SRB. In addition to clean speech, SRB also uses inherently noisy speech containing specific types of perturbations that are hard to simulate, but are likely to be encountered by ASR models in the real-world. Specifically, SRB uses accented speech from Common Voice [24], and inter-personal conversations and far-field recordings from CHiME-6 [7] and AMI [8].
3.2 Perturbations
The taxonomy of the perturbations used in our benchmark is presented in Fig. 2. The perturbations are of two broad types: (1) non-adversarial and (2) adversarial. The non-adversarial perturbations fall under one of four categories: (1.1) Inter-personal communication gathered in social settings using near-field and far-field microphones. This setting is representative of the environments in which voice-enabled smart devices are expected to operate. The data for this category are drawn from the ChiME [7] and AMI [8] corpora. (1.2) Environmental effects, such as white noise, environmental noise, spatial acoustics (room impulse response and echo) (1.3) Digital augmentations produced by standard audio processing operations, and special effects. The audio processing operations included amplitude gain, resampling, lowpass, and highpass filtering. The special effects included bass gain, treble gain, tempo increase, tempo decrease, speed increase, speed decrease, pitch increase, pitch decrease, chorus, tremolo, and phaser. All special effects are applied using SoX.  [^1] https://sourceforge.net/projects/sox/ (1.4) Speech Variations are represented by speech from non-American speakers, and speech generated by a TTS model [59]. Since users from a variety of ethnic backgrounds may interact with ASR models, they must be robust to a variety of accents. With the rise of generative AI, robustness to synthetic speech will also be important.
Figure 2: Taxonomy of perturbations currently used in SRB.
The adversarial perturbations we use are untargeted adversarial attacks whose objective is to prevent the correct transcription of input speech. We use two types of adversarial attacks: (2.1) utterance-specific and (2.2) utterance-agnostic attacks. The utterance-specific attack optimizes a perturbation for a given speech recording, such that a given model maximally mistranscribes. The utterance-agnostic attack works similarly but the perturbation is jointly optimized over many speech recordings. Thus, this perturbation can affect several speech recordings, even those that were not used during the optimization. This represents a more realistic scenario where a malicious actor tries to mount a denial-of-service style attack against an ASR model by introducing utterance-agnostic perturbation at some point in the transcription pipeline.
Most perturbations, except for those categorized as informal speech, or speech variations, are applied at 4 levels of severity, which are modulated by adjusting the perturbation parameters. Details about the perturbations and their parameters are presented in Appendix A.

### v3

@@ Abstract
As Automatic Speech Recognition (ASR) models become ever more pervasive, it is important to ensure that they make reliable predictions under corruptions present in the physical and digital world. We propose Speech Robust Bench (SRB), a comprehensive benchmark for evaluating the robustness of ASR models to diverse corruptions. SRB is composed of 114 challenging speech recognition scenarios which largely cover the range of corruptions that ASR models may encounter when deployed in the wild. We use SRB to evaluate the robustness of several state-of-the-art ASR models and observe that model size and certain modeling choices such as the use of discrete representations, or self-training appear to be conducive to robustness. We extend this analysis to measure the robustness of ASR models on data from various demographic subgroups, namely English and Spanish speakers, and males and females. Our results revealed noticeable disparities in the model’s robustness across subgroups. We believe that SRB will significantly facilitate future research towards robust ASR models, by making it easier to conduct comprehensive and comparable robustness evaluations.
@@ 1 Introduction
As novel ML models continue to be developed and deployed at an ever-increasing rate, it has become crucial to ensure their robustness to challenging real-world scenarios, where corruptions arising from a myriad of sources, including the environment, sensing apparatus, and even malicious actors are present. To this end, prior works have developed comprehensive robustness benchmarks, particularly for vision (Hendrycks & Dietterich, 2019; Hendrycks et al., 2021a; b; Croce et al., 2020) and natural language processing models (Wang et al., 2021a; 2022b), that evaluate a model’s performance under a variety of challenging scenarios. These benchmarks have proven to be invaluable to the advancement of research into more robust models because (1) they unify robustness evaluations, thus enabling meaningful comparisons across models and allowing progress to be accurately tracked, and (2) they make it easier for researchers to comprehensively evaluate the robustness of their models by aggregating a diverse and representative set of scenarios, and methods of simulating them, in a single benchmark.
@@
While several robustness benchmark datasets exist for Automatic Speech Recognition (ASR) models (Barker et al., 2017; Kraaij et al., 2005; Wichern et al., 2019; Reddy et al., 2020; Cosentino et al., 2020; Hershey et al., 2016; Chen et al., 2020; Snyder et al., 2015; Kinoshita et al., 2013; Ko et al., 2017; Nakamura et al., 2000; Jeub et al., 2009), none of the currently existing ones are in any sense comprehensive, because each benchmark measures the model’s robustness w.r.t. to one or a few specific types of corruptions or scenarios, which puts the onus on model developers to find and collect all the relevant benchmarks to evaluate their model comprehensively. This has often resulted in model developers evaluating their models on disparate benchmarks  (Radford et al., 2023; Wen et al., 2016; Chen et al., 2022; Likhomanenko et al., 2020), which makes it hard to reliably compare performance and robustness across models. Recently, Huggingface Open ASR Leaderboard (Srivastav et al., 2023) has sought to unify ASR model evaluations by developing a benchmark consisting of several real-world speech datasets. Although evaluating models on exclusively natural data may accurately reflect average case real-world performance, it is generally not informative about the specific types of corruptions the models are weak against, because the noise sources present in these datasets are not controlled or even fully known. For example, the crowdsourced recordings in Common Voice (Ardila et al., 2019) contain a variety of distortions including sensor noise from low-quality equipment, background noise, and mispronunciation by non-native speakers. Furthermore, digital perturbations like special effects, computer-generated speech, and adversarial examples, that may be prevalent in digital content are largely overlooked by existing benchmarks.
@@
In this paper, we propose Speech Robust Bench (SRB), a benchmark for comprehensively evaluating the robustness of ASR models to input perturbations and corruptions. SRB is designed to address the aforementioned major shortcomings of existing ASR robustness benchmarks, i.e., that (1) they are often specialized and thus are not individually comprehensive, (2) even taken together, they overlook important challenging scenarios, like special effects and adversarial attacks, and (3) may not reveal the specific weaknesses of the models. SRB addresses these shortcomings by evaluating ASR models under a comprehensive set of challenging scenarios, using recordings that are either are recorded under specific scenarios, and thus are inherently “noisy”, or recordings that are digitally perturbed to simulate the various scenarios. SRB uses real recordings of accented speech and inter-personal conversations to evaluate robustness to articulatory and lexical variability. We take care to ensure that the recordings are clean and do not have any other corruption that may confound the results. To digitally simulate challenging scenarios, we curate a large comprehensive bank of 114 perturbations that represent common distortions arising from the environment, recording equipment, special effects, computer-generated speech, and adversarial attacks that are often overlooked by existing benchmarks.
@@
To highlight the need for and the benefits of doing systematic and comprehensive robustness assessment, we evaluate the robustness of several popular ASR models (§4.1) using SRB. We observe that Whisper (Radford et al., 2023) is the most robust on average among the models we tested, even outperforming the more recent Canary (NVIDIA, ). We conduct further analyses to disentangle the effects of model and training data size, revealing that larger models tend to be more robust than smaller models, even if the latter are trained on significantly more data. We further extend our analysis by evaluating the models’ robustness for the various population sub-groups, namely, English and non-English (Spanish) speakers, and male and female speakers. We find that significant disparities exist across these sub-groups, thus identifying issues with fairness of the models and highlighting areas where future work could provide clear improvements. Besides pinpointing robustness issues, this demonstrates the utility of SRB for fairness evaluations that consider robustness disparities across models as well.
@@
  • To facilitate out-of-the-box robustness evaluations for the research community, we have publicly released a large dataset (link omitted) containing perturbed versions of LibriSpeech (Panayotov et al., 2015), TEDLIUM (Hernandez et al., 2018) and Spanish Multilingual LibriSpeech (Pratap et al., 2020) test sets, and accented speech from common voice and segemented near- and far-field audios from CHiME-6 (Reddy et al., 2020) and AMI(Kraaij et al., 2005).
  • We open source our code with clear documentation to allow for easy reproducibility and extensibility. (https://anonymous.4open.science/r/speech_robust_bench-4C77)
@@ 2.1 Robust Automatic Speech Recognition
Several techniques have been proposed for making Automatic Speech Recognition (ASR) models robust to input perturbations, such as noise and other signal corruptions (Li et al., 2014). We can divide these techniques into two high-level categories: i) model-based and ii) feature-based. Model-based techniques modify the models to make them more robust. Examples of such approaches include adapting pre-trained models (Yu et al., 2009; Juang & Rahim, 1996), de-noising the audio before processing (Mohammadiha et al., 2013; Wilson et al., 2008), and training ASR models on noisy data (Likhomanenko et al., 2020). Since model-based strategies generally require access to noisy data (Li et al., 2014), they are most effective if the sources of noise, and/or the exact environment in which the ASR model will be deployed in are known, and one can gather data to represent those. Feature-based approaches, on the other hand, involve developing handcrafted features, invariant to noise and corruptions in the signal (Li et al., 2014). Several of these features are inspired by biological audition (Kim & Stern, 2016; Hermansky et al., 1991; Hermansky & Sharma, 1998), while others use signal processing techniques (Li et al., 2014). Generally, these methods are designed to extract the components of the audio signal salient for speech production and perception, while discarding irrelevant components (Stern & Morgan, 2012). Consequently, they do not require precise knowledge of the environment and noise distributions. Recently, however, handcrafted features have fallen out of favor, and have been replaced by features learned via end-to-end training of deep learning models on large amounts of data (Baevski et al., 2020; Hsu et al., 2021a; Likhomanenko et al., 2020; Radford et al., 2023). Proponents of these techniques posit that models trained on larger datasets become more robust. Our evaluations in § 4 reveal that there are several input perturbations against which smaller models trained on less data outperform larger models trained on more data.
@@ 2.2 Adversarial Robustness
Adversarial perturbations can change the response of a model when added to their inputs, but are either imperceptible to humans or perceptually and semantically irrelevant enough to be ignored by them (Szegedy et al., 2014; Goodfellow et al., 2014). Adversarially perturbed inputs are known as adversarial attacks. They can be targeted (aiming to change a prediction to a specific incorrect class), or un-targeted (aiming to change a prediction to any incorrect class (Akhtar et al., 2021)). The design of adversarial attacks is determined by the level of knowledge the attacker is assumed to have about the target model. Attacks that assume full knowledge of the target model’s architecture and weights (white-box threat model) often use gradient-based optimization techniques (Szegedy et al., 2014; Goodfellow et al., 2014; Madry et al., 2018; Laidlaw et al., 2021; Akhtar et al., 2021). Attackers who do not have any knowledge of the target model’s architecture and only have query access to it (black-box threat model) typically use gradient-free optimization methods (Wang et al., 2022a; Andriushchenko et al., 2020; Wicker et al., 2018; Chen et al., 2017; Zhao et al., 2020; Vo et al., 2022). An intriguing property of adversarial perturbations is that they transfer between models (Papernot et al., 2016), and inputs (Akhtar et al., 2021; Neekhara et al., 2019), i.e. perturbations designed for one model/input may be effective against others as well. Our SRB includes two types of white box adversarial attacks; those that generate perturbations: 1) specific to each input (Madry et al., 2018), 2) that cause models to mis-transcribe multiple inputs (Neekhara et al., 2019).
@@ 2.3 Robustness Benchmarks for Speech
While several robustness benchmark datasets exist for ASR models, they, unfortunately, suffer from three major shortcomings that make it difficult to perform robustness evaluations in a way that is comprehensive, sufficiently fine-grained, and comparable across models.
@@
Firstly, we find that existing benchmarks do not comprehensively evaluate robustness. While past works such as (Pearce & Picone, 2002) did indeed propose benchmarks containing diverse perturbations, they were limited to relatively simple data unsuitable for modern ASR models. However, many recently proposed benchmarks measure the robustness in one or a few specific types of scenarios. For example, some datasets (Kinoshita et al., 2013; Nakamura et al., 2000; Jeub et al., 2009; Ko et al., 2017) focus exclusively on reverberant speech, while others focus only on multi-speaker scenarios (Kraaij et al., 2005; Barker et al., 2017), environmental noise (Snyder et al., 2015; Reddy et al., 2020; Wichern et al., 2019; Hirsch & Pearce, 2000), or accented speech (Lander, 2022; Shi et al., 2021). Furthermore, certain types of digital perturbations, like special effects, computer-generated speech, and adversarial examples, which humans are usually invariant to (and thus ASR models are expected to be as well) are largely overlooked by existing benchmarks. SRB evaluates robustness under a wide range of scenarios, including those represented in prior works, and novel ones that have not yet received due attention, thus enabling comprehensive robustness evaluations via a single benchmark.
Secondly, since there are several robustness benchmarks that researchers can choose from, the onus is on model developers to collect all relevant benchmarks and comprehensively evaluate their model. This has often resulted in model developers evaluating their models on disparate benchmarks  (Radford et al., 2023; Wen et al., 2016; Chen et al., 2022; Likhomanenko et al., 2020), which makes it hard to reliably compare performance and robustness across models. For example, several works have used subsets of existing benchmarks to evaluate the model robustness (Radford et al., 2023; Wen et al., 2016; Chen et al., 2022), or have evaluated the robustness of models by computing their transcription accuracy on multiple natural speech datasets (Likhomanenko et al., 2020; Radford et al., 2023; Hsu et al., 2021b). SRB covers a wide range of challenging ASR scenarios in a single benchmark, it obviates the need to select from various benchmarks, and, thus, can unify robustness evaluations across studies and make them comparable.
@@
Finally, robustness benchmarks that try to mitigate the above two shortcomings are often too coarse to reveal the specific scenarios and corruptions the model(s) struggle against. For example, Huggingface Open ASR Leaderboard(OAL, Srivastav et al. 2023) evaluates models on several real-world speech datasets, which may accurately reflect their average-case real-world performance, it may not inform about the specific type of perturbations the models are weak against. This is because noise sources present in these datasets are not controlled or even fully known. For example, the crowdsourced recordings in Common Voice (Ardila et al., 2019) contain a variety of distortions including sensor noise from low-quality equipment, background noise, and mispronunciation by non-native speakers. If a model struggles with Common Voice, identifying the specific types of noise it is sensitive to is not a straightforward task. In the design of SRB, we have ensured that the sources of noise (and variance, in general) in each utterance are limited and known, which allows users to pinpoint specific scenarios and/or perturbations under which their models struggle (e.g. Fig. 3 and § 4).
@@ 3 Speech Robust Bench
Speech Robust Bench (SRB) evaluates the robustness of ASR models by a three-step process consisting of (1) scenario simulation, (2) transcription, and (3) metrics computation, as shown in Fig. 1: first, various challenging speech recognition scenarios are simulated by applying a large bank of synthetic perturbations to clean speech datasets (§ 3.1), as well as by using inherently noisy speech datasets with limited and known sources of real noise and variations that are difficult to simulate. Next, the perturbed recordings, the original clean recordings, and the recordings with inherent noise are transcribed using the target ASR model. Finally, the predicted and reference transcripts are compared, and the accuracy and robustness of each model in each setting is captured with various metrics (§ 3.2). To account for the differences in the level of difficulty between scenarios, we also estimate speech quality scores using appropriate models and use them to calculate normalized metrics.
Figure 2: Taxonomy of scenarios currently represented in SRB. Scenarios in dashed boxes have real-world recordings, while scenarios in solid boxes are simulated by digitally adding perturbations.
3.1 Scenario Simulation
The various speech recognition scenarios simulated by SRB are taxonomized in Fig. 2, and can be divided into six high-level categories, namely (1) clean speech, (2) social gatherings, (3) speech variations, (4) environmental effects, (5) digital augmentations, and (6) adversarial attacks. The scenarios are described briefly below, while more details are given in Appendix A.
(1) Clean speech: SRB uses clean speech for two purposes: to benchmark the baseline accuracy of ASR models, and to simulate various challenging scenarios by perturbing it. Clean speech is drawn from Librispeech (Panayotov et al., 2015) test-clean, TEDLIUM (Hernandez et al., 2018) release 3 test and MultiLingual LibriSpeech (MLS) (Pratap et al., 2020) test. LibriSpeech contains professional recordings of English audio books. Meanwhile, TEDLIUM contains professional recordings of English TED talks and provides lexical and phonetic diversity which LibriSpeech may lack. To increase the applicability of SRB to non-English and multi-lingual models we also include Spanish speech from MLS, which contains professionally recorded audio books in several languages.
(2) Social Gatherings: The ability to transcribe speech from semi-formal or informal settings, as well as far-field audio is useful for models used in meeting rooms, smart homes, and even subtitle generation, thus SRB includes English speech from dinner parties and meetings recorded by (2.1) near- and (2.2) far-field mics from CHiME-6 (Barker et al., 2017) and AMI (Kraaij et al., 2005).
(3) Speech Variations: ASR models must remain accurate under variations in pronunciation and prosody to serve diverse speakers. We therefore include (3.1) clean accented speech [^1] Accent annotations (excluding US English), and a DNSMOS score $\geq 3.4$ (Reddy et al., 2020). from English and Spanish subsets of Common Voice 17 (CV17, Ardila et al. 2019) in SRB. To provide additional prosodic variability and to represent the increasing pervasiveness of generative AI, we also include synthetic speech generated by YourTTS (Casanova et al., 2022) (English) and Bark Suno (b) (Spanish) from transcripts of the three clean datasets from scenario (1) in the voices of English and Spanish speakers from VCTK (Yamagishi et al., 2019) and Bark Speaker Library (v2, Suno a), respectively.
The following scenarios involve synthetic perturbation of all three clean datasets from scenario (1).
(4) Environmental Effects: While noisy real speech datasets like CommonVoice (Ardila et al., 2019) and Switchboard (Godfrey et al., 1992) exist, the noise in them is not controlled or even known. Thus in SRB, we perturb clean speech to simulate (4.1) environmental noise, and (4.2) spatial acoustics. Concretely, we add real environmental noise from ESC-50 (Piczak, 2015), MS-SNSD (Reddy et al., 2019), MUSAN (Snyder et al., 2015) and WHAM! (Wichern et al., 2019) at Signal-to-Noise Ratios (SNR) of 10, 20, 30 and 40 dB. To simulate spatial acoustics, we add echo via SoX [^2] Available from https://sourceforge.net/projects/sox/. and simulate Room Impulse Response (RIR) via convolution with real and simulated RIRs from Ko et al. (2017).
(5) Digital Augmentations: Digital media often undergoes processing and contains special effects, which are therefore included in SRB. Specifically, we include standard audio processing operations like amplitude gain, resampling, lowpass, and highpass filtering, (2.2) special effects like bass gain, treble gain, tempo increase, tempo decrease, speed increase, speed decrease, pitch increase, pitch decrease, chorus, tremolo, and phaser, and (2.3) Gaussian white noise.
(6) Adversarial Attacks: Models used in high-stakes settings are prime targets for adversaries and thus must resist attempts to compromise their accuracy. We use two types of adversarial attacks in SRB: (2.1) utterance-specific and (2.2) utterance-agnostic attacks. The utterance-specific attack searches for a perturbation $\delta$, for a given speech recording $x$, such that a given model maximally mistranscribes it. To find $\delta$, we follow Madry et al. (2018) and use projected gradient descent to solve $\max_{\delta:\text{SNR}(\delta,x)\leq\epsilon}\mathcal{L}(M(x),y^{*})$, where $\mathcal{L}$ is a differentiable loss function, like CTC-Loss, between the model’s output $M(x)$ and the true transcript $y^{*}$, with $\epsilon\in[10,40]$. The utterance-agnostic attack is similar to the utterance-specific attack, except $\delta$ is optimized over a held-out set, $\mathcal{X}^{\text{dev}}$, instead of each test utterance. This represents a more realistic scenario where an attacker tries to mount a denial-of-service attack against an ASR model by introducing utterance-agnostic perturbation at some point in the transcription pipeline. We use the method of Neekhara et al. (2019) to find $\delta:\mathbbm{E}_{x\in\mathcal{X}^{\text{dev}}}\text{SNR}(\delta,x)\leq\epsilon\in[10,40]$ such that $CER(\{x+\delta|x\in\mathcal{X}^{\text{dev}}\})>\tau$ (see Alg. 1), where $\mathcal{X}^{\text{dev}}$ is the dev split of LibriSpeech, TEDLIUM and MLS.
Note of Extensibility and Usage: We have released our source code with instructions for reconstructing the data in SRB, and reproducing the results of this paper ( § 1). SRB can easily be extended to other languages and speech datasets using the provided scripts for extracting accented speech from any language in CV17, and for simulating scenarios 3.2-6 on any speech recording or dataset.
We have also publicly released the data for all the above scenarios, except adversarial attacks and perturbed TEDLIUM recordings, on Huggingface Hub [URL omitted for double-blind]. We made these exceptions because TEDLIUM’s license prohibits the distribution of derivatives, and the adversarial attacks must be computed separately for each target ASR model. To the extent possible, we encourage users to evaluate their models on the publicly released data to ensure reproducibility.